### v3

# Eaman Jahani

Abstract
The recent surge in generative AI has led to new models being introduced almost every month. In light of this rapid progression, we pose and address a central question: to what extent must prompts evolve as the capabilities of generative AI models advance? To answer this question, we conducted an online experiment with N = 1,893 participants where each participant was incentivized to write prompts to reproduce a target image as closely as possible in 10 consecutive tries. Each participant was randomly and blindly assigned to use one of three text-to-image diffusion models: DALL-E 2, its more advanced successor, DALL-E 3, or a version of DALL-E 3 with automatic prompt revision. In total, we collected and analyzed over 18,000 prompts and over 300,000 images. We find that task performance was higher for participants using DALL-E 3 than for those using DALL-E 2. This performance gap corresponds to a noticeable difference in the similarity of participants’ images to their target images, and was caused in equal measure by: (1) the increased technical capabilities of DALL-E 3, and (2) endogenous changes in participants’ prompting in response to these increased capabilities. Furthermore, while participants assigned to DALL-E 3 with prompt revision still outperformed those assigned to DALL-E 2, automatic prompt revision reduced the benefits of using DALL-E 3 by 58%. Our results suggest that for generative AI to realize its full impact on the global economy, people, firms, and institutions will need to update their prompts in response to new models. Not doing so could leave more than half of the potential benefits of these AI systems untapped.
Recent economic forecasts predict that generative AI could add trillions of dollars to the global economy annually, and there is mounting evidence that this new technology is being integrated into work practices in areas as diverse as business, medicine, and government FactSet (2024); Zhang and Kamel Boulos (2023); Bright et al. (2024). This increasing generative AI adoption appears to be warranted, with recent studies finding that the adoption of generative AI improves the productivity of workers across various professional settings (Brynjolfsson et al., 2023; Dell’Acqua et al., 2023; Noy and Zhang, 2023). A common feature of these studies is that the model participants interacted with did not change; however, generative models are continuously updated, with new versions being released as often as once a month. Humans interface with generative models by providing textual instructions or “prompts” to the models (Oppenlaender, 2023; Schulhoff et al., 2024; Don-Yehiya et al., 2023; Xie et al., 2023). It remains an open question whether organizations will need to adapt these prompts as models improve to realize the full economic and societal benefits of generative AI. For example, it is now common amongst software companies to have a hundred or more development teams, each with dozens, if not hundreds, of prompts integrated with their code. On the one hand, firms might write these prompts once and subsequently benefit from new models with minimal additional effort. Alternatively, as models improve, organizations may need to continually adapt their prompts. In this paper, we investigate whether this prompt adaptation process is necessary, and the extent to which the ability to adapt one’s prompts is a specialized skill.
We do so by conducting a pre-registered online lab experiment in which $N=1,893{}$ participants from Prolific were randomly and blindly assigned to complete the same task using one of three generative AI models with varying capabilities: (1) DALL-E 2, (2) its more advanced successor DALL-E 3, or (3) a version of DALL-E 3 with automatic large language model (LLM)-based prompt revisions (hereafter referred to as “DALL-E 3 with revision”). [^1] The main text presents a subset of pre-registered analyses, with additional results and deviations from pre-registration detailed in the SI. The LLM-based prompt revisions in the third condition were not made visible to participants during the experiment.222prompt revision is the default behavior of the DALL-E 3 API endpoint. We followed the DALL-E 3 API documentation and attempted to disable prompt revision in the second, “DALL-E 3” treatment arm by prepending a system prompt to participants’ prompts. While this drastically reduced the number and extent of LLM-based prompt revisions, it did not completely eliminate them. Therefore, the “DALL-E 3” treatment arm still includes some automatic prompt revisions. For brevity and simplicity, we refer to this treatment arm as “DALL-E 3” throughout the paper. Full details on the system prompt’s efficacy are provided in the SI. The task required each participant to make at least 10 attempts at recreating a “target image” as closely as possible by prompting their assigned model. Each participant’s target image was randomly selected from a curated set of 15 images; this set of images was constructed to span realistic use cases of AI-generated images. Throughout the task, participants could view all of their past prompts and generated images alongside the target image. Participants were paid \$4 USD to complete the task, and to incentivize performance, they received \$8 USD if they were in the top 20% of participants, as measured by the similarity of their best attempt to the target image.
Results
To understand how organizations might benefit from model improvements, we first examined whether access to a more advanced generative model enhanced task performance. Comparing the DALL-E 2 and DALL-E 3 treatment arms, we find that providing access to a more advanced generative model improved task performance. Participants using DALL-E 3 produced images that were, on average, $z=0.19$ standard deviations closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$), and the size of this gap did not shrink as participants gained more experience with the task (See Figure 1B; $\beta=0.0010$, $p=0.0227$, see SI for full regression). Figure 1A presents three examples from our data to illustrate the qualitative magnitude of this average treatment effect (ATE), which corresponds to a considerable increase in qualitative similarity to the target image.
[…]
Access to a more advanced generative model not only increased task performance but also induced changes in the way that participants wrote prompts, which we highlight through two exploratory analyses. First, prompts to the more advanced model were longer and contained more information. Those assigned to DALL-E 3 wrote prompts that were 24% longer ($\Delta Words=6.9$, $p<10^{-5}$), and this difference in prompt length increased as participants made successive replication attempts (Figure 1C; $\beta=1.17$ extra words per attempt, $p<10^{-8}$). Despite this difference in length, prompts written for both models contained similar proportions of nouns and adjectives (48% DALL-E 3 vs. 49% DALL-E 2; $p$ = $0.215$), suggesting that prompts submitted to DALL-E 3 conveyed additional descriptive information. Second, those assigned to DALL-E 3 explored the space of possible prompts differently; they wrote prompts that were more similar to one another, both sequentially from prompt-to-prompt (as shown by the color scale in Figure 1C; $\beta=0.0184$, $p=0.02$) and in the aggregate ($\beta=0.0191$, $p=0.008$) (see the SI for the details of our prompt-related textual analyses). In other words, those assigned to DALL-E 3 were more likely to thoroughly explore a smaller set of prompting approaches.
The changes in prompting we observe are an important mechanism through which participant performance improves, suggesting that organizations will need to adapt their prompts as models evolve. More specifically, using the decomposition procedure described in the Materials and Methods, we find that changes in prompting accounted for 48% of the performance ATE ($\Delta CoSim=0.00788$, $p=0.024$; blue in Figure 2), whereas the shift to a more capable model accounted for 51% of the performance ATE ($\Delta CoSim=0.00841$, $p<10^{-8}$; red in Figure 2).
We also supplied the prompts written for DALL-E 3 to DALL-E 2 (see Materials and Methods for details). Doing so did not improve performance ($\Delta CoSim=0.0020$, $p=0.56$), indicating that DALL-E 3 prompts specifically took advantage of its superior ability to render information (Betker et al., 2023).
Figure 2: The decomposition of the ATE (black) into model (red) and prompting (blue) effects. The x-axis indicates the model for which the prompts were written (above the dashed line) and the model used for replay (below the dashed line). Effects are shown relative to a baseline of prompts written for and replayed on DALL-E 2. Error bars show 95% confidence intervals based on bootstrapped standard errors clustered by participant.
Our results show that adapting prompts is crucial for maximizing the benefits of more capable generative models, raising the question of how organizations should approach this adaptation process. One possibility is that LLM-based, automated prompt revision (Li et al., 2024; Betker et al., 2023) obviates the need for humans to tailor their prompts to different generative models. Comparison of our third treatment arm–DALL-E 3 with revision–to the DALL-E 3 arm highlights the potential pitfalls of this approach. We find that prompt revision actually caused participants to perform worse in our context as it reduced the benefit of using DALL-E 3 by nearly 58% (95% CI: [40%, 76%]).
Furthermore, the positive impact of access to DALL-E 3 with revision performed only slightly better than those assigned to DALL-E 2 ($\Delta CoSim=0.0069$; $p=0.042$) and was less than the positive impact of directly passing prompts written for DALL-E 2 to DALL-E 3 (i.e., the model effect shown in Figure 2). These findings suggest that as currently implemented, AI-assisted prompt revisions are not a panacea. They can actually inhibit people’s capacity to leverage a model’s capabilities when misaligned with an end user’s goals.
[…]
Discussion
The primary limitation of our work is that we only studied the transition from DALL-E 2 to DALL-E 3. We leave the study of future transitions and different types of generative AI models to future work. Should our results generalize to other models, then we would predict that firms that do not update their prompts as models progress are leaving roughly half the benefit of AI unrealized. Software systems, models, and prompts are becoming increasingly interwoven which is a trend we expect to accelerate. Thus, the lack of prompt maintenance could, in turn, dampen the economic impact of AI on the global economy. Fortunately, updating prompts can be learned by humans across the skill distribution even if models are not currently able to update prompts in an automated fashion. Thus, we anticipate that as firms adopt generative AI, they will need to regularly refine prompts to fully realize the potential of evolving models. In particular, one could imagine a lock-step dynamic where, as models continually improve, people will need to—and are easily able to—respond by adapting their prompts to take advantage of the newest model’s capabilities. Such a pattern suggests that as generative AI models advance, technical prompting infrastructure will not be a one-time investment. Rather, prompting will be the mechanism by which people and firms unlock new models’ capabilities.
Table 1: Model and Prompting Effect by Performance
Effect Estimate p-value
Model $0.011000$ $p<0.00001$
$(0.001860)$
Model $\times$ $-0.000060$ $0.0210$
Performance Decile $(0.000030)$
Prompting $0.010900$ $0.0006$
$(0.003160)$
Prompting $\times$ $-0.000060$ $0.2444$
Performance Decile $(0.000050)$
Material and Methods
Outcome measure Given the stochastic nature of generative AI models, our outcome measure of task performance is the expected similarity of images generated from each prompt to the corresponding target image. To compute this, for each prompt we generated 10 images and computed CLIP embeddings (Radford et al., 2021) for both these and the target images. Each prompt’s expected similarity score was then computed as the mean cosine similarity ($CoSim$) between its generated image embeddings and the target image embedding (see SI for details).
Effect Decomposition We conducted an exploratory analysis where we regenerated images with all participant prompts on both models. This allows us to decompose the ATE into two components: the “model effect” (the average improvement when running DALL-E 2 prompts through DALL-E 3) and the “prompting effect” (the difference in target similarity between DALL-E 3 and DALL-E 2 participants’ prompts, both evaluated on DALL-E 3). Note that all main text analyses use “replay” data from Figure 2 for consistency and to avoid model drift issues. All results hold with the “original” data. See the SI for details.
Experimental Materials Preregistration, data, and analysis code will be deposited at https://osf.io/ejbtp. Full details regarding our experiment design, analysis techniques, and additional results can be found in the SI.
Supplementary Information (SI)
Appendix A Experiment Design
A.1 Task Design
Participants were asked to reproduce a single target image as closely as possible using a text-to-image generative AI model (i.e., DALL-E 2, DALL-E 3 without prompt revision, or DALL-E 3 with prompt revision, all developed by OpenAI). They did so by successively submitting prompts. In response to each submitted prompt, the model would generate an image, which was then displayed to the participant next to their assigned target image. Participants were instructed to make at least 10 attempts at trying to recreate the target image within a 25-minute window, with no upper limit on their number of attempts.
All interactions between participants and the generative AI models occurred on a custom-built online interface designed to resemble OpenAI’s ChatGPT interface but with some adjustments related to our task (e.g., displaying the target image and the total number of attempts so far to the user). On the right-hand side of the interface, participants were shown the target image they were randomly assigned to recreate. On the left-hand side, participants were shown their previously submitted prompts as well as the resulting generated images. We placed the text box where participants were able to write and submit their prompts at the bottom of the interface. Prompts were limited to a maximum of 1,000 characters. Participants were informed that their interactions with their assigned model would be memory-less, i.e., the model retained no memory of previous prompts and only used the current prompt to generate each image. Before the task, participants were provided with written and video instructions on how to interact with our experiment interface. Our task did not assume nor require prior experience with any generative AI tools.
After the task, we surveyed participants’ opinions and preferences regarding generative AI tools. We also inquired about their self-assessed occupational skills and how often they 1) engaged in creative writing, 2) wrote specific instructions, and 3) engaged in any sort of computer programming. Finally, we collected socio-demographic data, such as age, gender, and occupation.
A.2 Randomization
We randomized participants across two dimensions: the target image and the text-to-image generative AI model that participants had access to. We randomized participants across both dimensions simultaneously using complete randomization, generating 45 possible target image-model cells. We conducted a balance check after the conclusion of the experiment with a $\chi^{2}$ test across all cells. With $\chi^{2}=7.056,\;df=44$, the resulting p-value equals to 1 and thus we cannot reject the null hypothesis that the proportions are equal across all 45 groups:
$H_{0}:p_{1}=p_{2}=\cdots=p_{45}$
Participants were unaware of this randomization.
A.2.1 Generative Models
We randomly assigned participants to 1 of 3 generative models:
  1. 1.
    DALL-E 2, which is referred to at points in the main text as the inferior model.
  2. 2.
    DALL-E 3 (Verbatim), which is referred to at points in the main text as the superior model.
  3. 3.
    DALL-E 3 (Revised), which is referred to at points in the main text as DALL-E 3 with Revision
Both the “verbatim” and “revised” versions of the DALL-E 3 treatment utilize the same underlying image-generating model; the distinction lies in the pre-processing applied before submitting user prompts to OpenAI’s image-generating API. OpenAI’s DALL-E 3 system, by design, employs a GPT-4 model to rewrite user prompts, adding more detail before processing the modified prompt using the DALL-E image-generating model. During our experiment, it was not possible to explicitly disable this prompt rewriting feature of the DALL-E 3 system. To manage this behavior, we defined two treatments utilizing the DALL-E 3 model.
In the DALL-E 3 (Revised) treatment arm, we submit the participant’s prompt directly to OpenAI’s API and do not interfere with the default prompt rewriting process. In the DALL-E 3 (Verbatim) treatment arm, we prepend a string instructing the GPT-4 model to not modify the participant’s prompt before passing it forward to DALL-E 3. This string is never visible to participants and was modeled after a prefix specifically suggested in OpenAI’s online documentation for the DALL-E 3 endpoint. [^3] See here and here for online documentation. We modified the recommended prefix slightly to account for the fact that we did not expect our participants to always submit “extremely simple prompts.” The string we prepended to prompts is found below:
“I NEED to test how the tool works with my prompt as it is written. DO NOT add any detail; just use it AS IS:”
Prepending this string to participants’ prompt did reduce the rate at which OpenAI’s endpoint modified prompts, but compliance was not perfect. Thus, we view the “verbatim” treatment arm as more of an intent-to-treat intervention. The GPT model still modified 59% of participant prompts. The average token sort ratio (TSR) between the original prompt and the modified prompt was 77 for the DALL-E 3 (Verbatim) arm, compared to an average token sort ratio of 44 across the entire DALL-E 3 (Revised) treatment arm (a TSR of 100 denotes an exact string match). Conditional on any modification (any observations with TSR $<$ 100), the average TSR between the original prompt and the modified prompt was 61 for the DALL-E 3 (Verbatim) arm, compared to an average TSR of 44 across the entire DALL-E 3 (Revised) treatment arm.
A.2.2 Target Images
We randomly assigned participants to 1 of 15 target images. The set of target images consisted of 5 images each from 3 different broad categories: business and marketing, graphic design, and architectural photography. We chose these to represent the use cases suggested by the prompt categories on https://promptbase.com/, a leading marketplace for image generation prompts. The images vary in color, style, content, and complexity within and across categories. These images can be found online linked to the pre-registration document: https://osf.io/ejbtp. As we discuss in Section F, performance, and variability of performance varied substantially across images. In other words, some images were much easier than others to replicate with the generative models, which we view as additional evidence that the set of 15 images was reasonably diverse.
A.3 Subjects
Our Prolific-recruited US sample (N = 2,059) was limited to fluent English speakers, and we prevented participants from completing the task more than once. We also prevented users from completing the task on mobile devices or tablets. Data was collected between December 12, 2023 and December 19, 2023. Participants were guaranteed a payment of \$4 USD for completing the task and could earn an additional \$8 USD (a 200% bonus) if they ranked in the top 20% of participants in DreamSim of their image most similar to the target (construction of DreamSim is described in section B.5.1). The median time to complete our entire task, including a demographic survey, was 22 minutes. Given that 20% of subjects received a bonus, the average compensation for participants in our study was \$5.60 USD per person, or about \$15 USD per hour. We explained the payment and incentive scheme to participants in full multiple times during the onboarding phase of the experiment, and asked participants to confirm their understanding before they were allowed to complete the task. The onboarding process also included multiple attention checks; participants who failed the first check were immediately disqualified. For subsequent checks, participants were required to retry until they demonstrated understanding.
A.4 Model Endpoints
We used the following model endpoints and parameters to generate images from prompts:
  1. 1.
    OpenAI API: We used the image generation endpoint of the official OpenAI Node.js library to generate images for user prompts during the experiment. For all treatment arms, we set the image size parameter to be 1024 x 1024 pixels. For the DALL-E 3 (Revised) and DALL-E 3 (Verbatim) treatment arms, we set the quality parameter to standard and the style parameter to natural.
  2. 2.
    Azure OpenAI Service: We used the image generation endpoints in the Python implementation of Azure OpenAI Service to generate all replay images based off user prompts collected during the experiment. For prompts replayed through the DALL-E 2 treatment arm, we deployed a set of DALL-E 2 models on Azure OpenAI Service and set the API version for each to the 2023-06-01-preview version. For DALL-E 2, we created replay images in batches of 5. For prompts replayed through the DALL-E 3 (Revised) and DALL-E 3 (Verbatim) treatment arms, we deployed a set of DALL-E 3 models on Azure OpenAI Service and set the API version for each to the 2023-12-01-preview version. The parameter values for image size, and quality and style for the DALL-E 3 treatment arms, were set to the same values as in the experiment.
Appendix B Measurement and Variables
B.1 Survey Data
For each participant, we used a Qualtrics survey to collect demographic information, information on the participant’s skills that may be relevant to generative AI use, and information on the participant’s attitude towards generative AI. This data includes:
  • Demographics: Ethnicity, Gender, Age, Highest level of education attained (some high school, high school, some college, associate’s degree, bachelor’s degree, master’s degree, doctoral degree, professional degree, other), Years of work experience, Annual Income (0-\$25k, \$25.001k-\$50k, \$50.001k-75k, \$75.001k-\$100k, \$100.001k-\$150k, \$150k+), and elicitation of sets of O*NET job skills that participants used in their occupation (reading comprehension, active listening, writing, speaking, critical thinking, social perceptiveness, coordination, instructing, programming, judgment and decision making, systems evaluations, science, active learning, learning strategies, monitoring, complex problem analysis, technology design, troubleshooting, quality control analysis, systems analysis).
  • Opinions and Skills: Computer programming proficiency and usage frequency (self-reported), Structured and creative writing proficiency and usage frequency (self-reported), Generative AI tool proficiency and usage frequency (self-reported), Attitudes towards net social impact of Generative AI (self-reported), Advice for (hypothetical) future participants on how to perform well on the task.
B.2 Prompt Data
For each prompt, we record the text of the participant’s prompt, the order in which it was submitted, the timestamp of submission, and for the DALL-E 3 treatment arms, the revised prompt returned by the model.
B.3 Image Data
For each prompt, the following images were collected:
  1. 1.
    The participant-facing images (OpenAI API endpoint): The image shown to the participant during the experiment, generated by the model they were assigned to using the prompt they submitted. These images were generated from December 12-19, 2023.
  2. 2.
    Post-hoc resampled images (Azure OpenAI endpoint): For any given prompt, the output of the text-to-image model is stochastic. To better approximate the expected image from a given prompt, we generated 20 additional images for each prompt after the experiment concluded. We provide full details on this procedure in Section C. These images were generated from December 26, 2023 - January 27, 2024. These images are not used for analyses presented in the main text, but were used for other pre-registered analyses. These additional analyses are discussed in Section F.
  3. 3.
    Post-hoc replayed images (Azure OpenAI endpoint): To decompose our overall effects into model and prompting effects, we generated “counterfactual images” for each prompt written under the DALL-E 2 and DALL-E 3 (Verbatim) treatments. In other words, we submitted all prompts written under both the DALL-E 2 and DALL-E 3 (Verbatim) treatments to both the DALL-E 2 and DALL-E 3 (Verbatim) endpoints. Similarly to the resampling procedure outlined above, we generated 10 images per prompt per model: we generated a single replay for each prompt-model pair from March 16-18, 2024, and then, to increase power, generated the replications for these replay images from June 14-27, 2024. This replay process produced a total of 20 images per prompt—10 under the original model, 10 under the counterfactual model. We re-submitted prompts to their original model to account for potential model drift, as this exploratory analysis was conducted multiple months after our initial data collection. For consistency, this replay data is used throughout the main text of our paper.
B.4 Sample Construction
The sample we analyze in the main text of our paper is constructed using the process described as follows.
  • The initial “raw” dataset collected during the experiment is comprised of 24,672 rows of raw prompt data (one prompt per row) generated by 2,059 participants.
  • We first removed rows with blank prompt entries, invalid prolific IDs, and unsuccessful attempts (logging errors). These exclusion criteria were pre-registered. This left us with 2,029 participants and 24,123 prompts.
  • We next removed participants from our sample if they failed to submit at least 10 prompts or if a participant submitted the same prompt at least five times in a row at any point during the task. Both of these exclusion criteria were pre-registered. These exclusion criteria were also explained to participants, who were told that payment was contingent on submitting at least 10 successful prompts and a “good-faith effort.” To avoid reward hacking, we did not specify the “no more than 5 repeated prompts” criterion for “good-faith effort.” This left us with 1,899 participants.
[…]
  • We next removed participants who failed to complete the Qualtrics survey. This exclusion criteria was pre-registered. This left us with 1,893 participants and 18,930 prompts.
  • We also removed prompts from our dataset according to a number of post-hoc, non-pre-registered exclusion criteria to ensure data quality and avoid selection bias. If a prompt had any of the following flags, it was removed from the sample:
    • Prompts sometimes trigger errors in OpenAI’s safety system because they contain language that might be deemed unsafe under OpenAI’s policies. The specific language that triggers these errors is constantly changing and not available publicly. If a prompt triggered a safety error during the replication or replay process, we re-submitted the prompt up to 50 times or until the 10 original arm replications/replay samples had been collected. We removed prompts if they failed to generate 10 replications on the original model or 10 replay samples under the counterfactual model during the replication/replay process. This affected 305 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms. It did not affect any DALL-E 3 (Revised) prompts, as we did not conduct replay analysis with the prompts from this treatment arm.
    • Due to rare latency issues, some prompts were assigned duplicate attempt numbers by the MongoDB database that we used to collect our data. This data collection error led to issues in the data analysis process. Thus, we excluded prompts with duplicate attempt numbers. This affected 34 prompts across all three treatment arms, and 20 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms, approximately 0.1% of the original data.
  • Our final sample included 1,893 participants and 18,560 prompts.
    B.4.1 “Off-Topic” Robustness Check
    While analyzing our data, we found that our sample contained a number of “off-topic” prompts that did not seem related to the task. As a robustness check on our main results, we used the following process to systematically identify and remove “off-topic” prompts. First, we generated embeddings for each prompt using OpenAI’s text-embedding-3-small model. We then calculated the mean embedding for each target image. Next, we calculated the Euclidean distance between each prompt’s embedding vector and the mean embedding vector for prompts corresponding to the focal prompt’s assigned target image. Finally, we removed the 2.5% of prompts that were most distant from the mean image-level prompt embedding vector. This led to the removal of 481 prompts across all three treatment arms, and 338 prompts between the DALL-E 2 and DALL-E 3 (Verbatim) treatment arms. All of our main text results are robust to the exclusion of these “off-topic” prompts.
B.5 Dependent Variables
B.5.1 Image Similarity
We pre-registered two quantitative measures of image similarity: the cosine similarity of CLIP embedding vectors and a recently developed measure called ‘DreamSim’ (Fu et al., 2023). In the main text, we present analyses using CLIP embedding cosine similarity, since it is likely more familiar to readers. Our results are qualitatively and quantitatively similar using DreamSim instead.
  • CLIP Embedding Cosine Similarity: To calculate CLIP embedding cosine similarity, we first generated CLIP embedding vectors (Radford et al., 2021) from Hugging Face (CLIP-embeddings) for each participant-generated image and for each target image. Unlike traditional image embeddings that only encode visual features, CLIP embeddings also capture semantic relationships between images and descriptive text. We then calculated the cosine similarity between each participant-generated image’s CLIP embedding and the relevant target image’s CLIP embedding.
  • DreamSim: DreamSim is an image similarity measure proposed recently by (Fu et al., 2023). The authors claim that relative to a measure such as CLIP embedding cosine similarity, DreamSim measures image similarity in a way that more effectively captures human visual perceptions of similarity. Because the original DreamSim metric outputs a distance measure, we invert this score $\tilde{D}=1-(\text{original DreamSim})$ to recast it as a similarity score. After doing so, both the inverted DreamSim and CLIP embedding cosine similarity are closer to 1 when two images are more similar and closer to 0 when two images are more dissimilar.
We find that these two measures of image similarity are highly correlated in our sample ($\rho_{pearson}=0.763$, 95% CI: [0.755 0.770]), and our main results are robust to the use of either measure. We present the results obtained when conducting our main text analyses using DreamSim in Section E.E.1.
B.5.2 Prompt Length
We measure the lengths of prompts written by participants in our sample, both in terms of the number of words in a given prompt and in terms of the number of characters in a given prompt. In our main text analysis, we present results only in terms of the number of words, since the two outcomes are highly correlated ($\rho_{pearson}=0.9954$, 95% CI: [0.99528, 0.99560]).
B.5.3 Embedding-based Prompt Similarity
We calculate two measures of embedding-based prompt similarity: successive similarity and aggregate similarity. Both measures use the vector embedding representation of each prompt in our sample, which we obtained using OpenAI’s text-embedding-3-small model (Neelakantan et al., 2022). The two similarity measures are defined as follows:
  • Successive similarity: The successive similarity ($ss$) is a measure of the similarity of a participant’s prompt to their immediately preceding prompt. We define the successive similarity of a prompt $p_{i,n}$ written by user $i$ to the their immediately preceding prompt $p_{i,n-1}$ as:
    $ss_{i,n,n-1}=\frac{\mathbf{E(p_{i,n})}\cdot\mathbf{E(p_{i,n-1})}}{||\mathbf{E(p_{i,n})}||\,||\mathbf{E(p_{i,n-1})}||},$ (1)
[…]
  • Aggregate similarity: The aggregate similarity ($as$) is a measure of how dispersed each user’s prompts are around their “average prompt” (calculated by taking the element-wise average of all prompt embeddings produced by the user). We define the aggregate similarity for the 10 prompts written by a given user as:
    $as_{i}=\frac{1}{10}\sum_{n=1}^{10}\lVert\mathbf{E(p_{i,n})}-\overline{\mathbf{E(p_{i,n})}}\rVert_{2}^{2},$ (2)
    where $\mathbf{E(p_{i,n})}$ is again the vector embedding representation of participant $i$’s $n^{\text{th}}$ prompt, $p_{i,n}$, and $\overline{\mathbf{E(p_{i,n})}}$ is the element-wise mean of all 10 of participant $i$’s prompts.
B.5.4 Successive Prompt Token Sort Ratio
Starting with each participant’s second prompt, we also calculated the token sort ratio (TSR) of each prompt $p_{i,n}$ to the immediately preceding prompt $p_{i,n-1}$. TSR is a fuzzy string-matching technique (Singla and Garg, 2012) that provides a continuous measure of how similar two strings are. We refer the reader to (Bosker, 2021) for a more in-depth description of how TSR is calculated.
B.5.5 Successive Prompt ‘Contains Previous Prompt’ Dummy
Starting with each participant’s second prompt, we record whether each prompt $p_{i,n}$ contains the immediately preceding prompt $p_{i,n-1}$ as an exact substring.
B.5.6 Prompt Composition
We use the spaCy v3.7.4 Python package’s en_core_web_sm model to tag the parts of speech (POS) in each prompt. SpaCy’s models utilize the ”universal POS tags” from the Universal Dependencies framework for grammar annotation Universal POS tags (2024). These tags encompass parts of speech such as adjectives, adverbs, nouns, and verbs. The model tags each word in a prompt according to this framework, after which we count the total number of words corresponding to each part of speech for each prompt.
B.5.7 Strategic Shifts
In addition to calculating the successive and aggregate similarity of prompts written by particular users, we also attempt to identify particular moments when participants shift their approach to prompting. In order to do so, we adapt a method proposed in (Torricelli et al., 2023) (because they are conducting research in a different context, (Torricelli et al., 2023) refer to these shifts as “topical transitions” as opposed to “strategic shifts”). To identify these strategic shifts, we first calculate the mean cosine similarity ($MCS$) for the embedding vectors of every possible pair of prompts submitted in response to a given target image, $t$:
$MCS_{t}=\frac{2}{P_{t}(P_{t}-1)}\sum_{a=1}^{P_{t}}\sum_{b=a+1}^{P_{t}-1}\textrm{CosineSim}(\mathbf{E(p_{a,t})},\mathbf{E(p_{b,t})}).$ (3)
where $P_{t}$ is the total number of prompts submitted in response to a given target image, and $a$ and $b$ are indices representing individual prompts for that target.
We then label any given prompt as a strategic shift ($SS$) if the cosine similarity of its embedding vector with that of the previous prompt is lower than this target-image-level mean:
$SS(\mathbf{p}_{i,n,t})=\begin{cases}1&\text{if }\textrm{CosineSim}(\mathbf{E(p_{i,n,t})},\mathbf{E(p_{i-1,n,t})})<MCS_{t}\\ 0&\text{otherwise}\end{cases}$ (4)
It is worth noting that Torricelli et al. (2023) uses the participant-level mean, as opposed to the task-level mean, as the cutoff for a topical shift. We instead use the task-level mean because in our setting, as it did not seem appropriate that half of each participant’s submitted prompts would be strategic shifts.
Appendix C Methods
C.1 Stratification
The results shown in the main text and in the SI are mostly stratified by reference image and iteration. In some analyses, we have stratified only on the reference image (e.g., for analyses presented at the iteration level). The exact stratification for each finding is indicated in section D. To stratify our results, we take a weighted average across $j=1,...,J$ cells defined by our stratification variables.
$\overline{Y}_{strat}=\sum_{j=1}^{J}\frac{N_{j}}{N}\bar{Y}_{j}$
To calculate the variance (and standard error) of this sample mean we apply the following:
$\widehat{\mathrm{Var}}(\overline{Y}_{strat})=\widehat{\mathrm{Var}}\left(\sum_{j=1}^{J}\frac{N_{j}}{N}\bar{Y}_{j}\right)=\sum_{j=1}^{J}\left(\frac{N_{j}}{N}\right)^{2}\frac{s_{j}^{2}}{N_{j}}$
where:
  • $N_{j}$ is the population size of stratum $j$.
  • $N$ is the total population size across all strata.
  • $\bar{Y}_{j}$ is the sample mean for stratum $j$.
  • $J$ is the total number of strata.
  • $s_{j}$ is the sample standard deviation of stratum $j$. Therefore, $s_{j}^{2}$ is sample variance stratum $j$.
C.2 Z-Scoring
We find statistically significant evidence for differences in the variability of performance across the 15 target images used in our experiments, which we discuss in Section F. In the main text, we also showed that performance increases across the attempts. To test whether our results are in some way due to this image-level or attempt-level variation, we replicate all analyses using the within-image-attempt Z-score of CLIP-cosine similarity of each image produced by participants in our experiment. Formally, this is:
$Z(\textrm{CosineSim}_{i,n,t})=\frac{\textrm{CosineSim}_{i,n,t}-\textrm{Mean}_{n,t}\left(\textrm{CosineSim}_{i,n,t}\right)}{\textrm{SD}_{n,t}\left(\textrm{CosineSim}_{i,n,t}\right)},$ (5)
where $\textrm{CosineSim}_{i,n,t}$ is the cosine similarity of user $i$’s image in attempt $n$ to target image $t$. The mean and standard deviation are computed per each image-attempt, but over the DALL-E 2 and DALL-E 3 treatment arms. We also use the rescaling above to test the robustness of our DreamSim-based analyses. Almost all the robustness analysis reported here use the Z-score scaled measure of performance within each image-attempt set. The only exception is any analysis that examines improvements and prompts across attempts, e.g. figure 1B, where Z-scores are computed only within each target image, and across all attempts of that image.
C.3 Accounting for Model Stochasticity
The output that generative AI models return in response to a given prompt is stochastic. The strength of this stochasticity is controlled by a model parameter referred to as the temperature, which could not be edited using the DALL-E API at the time of our experiment. To account for this model stochasticity, we generated 10 images for each prompt submitted by participants for all arms. We were then able to calculate the similarity between each replication and its corresponding target image, and calculate an “expected” CLIP cosine similarity and DreamSim score for each prompt by averaging over these samples. Given the replicated images, we can also calculate the standard deviation of cosine similarity induced by this stochasticity. With the expected cosine similarity and its standard deviation per prompt, we can compute a normalized Z-score for the observed image relative to its replication distribution. This Z-score measures the extent to which the observed image is better or worse than what’s expected for that prompt and will be used for further analysis in section F.
We generated these additional samples for both the original prompts on their assigned treatment arms, as well as replaying on the counterfactual arms, as introduced in Figure 2 in the main text. Importantly, OpenAI updated its content filters between our initial experiment and image re-sampling. As a result, some prompts that originally produced images either generated no images or fewer images than requested during our regeneration attempts. This affected 1.8% (371 out of 18,990 prompts) of the data in our sample under the ”replaying” procedure (Section B.B.3.3).
Appendix D Main Text Analyses
D.1 Task Performance and ATEs
The top pane of Figure 1B compares the average performance across models and attempt numbers (also referred to as iterations). It shows the average cosine similarity score stratified by the reference image. A notable feature in this figure is the performance dip during the second recreation attempt across both treatment arms. This is likely due to participants’ initial misunderstanding of the model’s “memoryless” nature. Participants failed to recognize that context from previous prompts was not carried over to new iterations. We observed numerous prompts in the second iteration across users that explicitly referenced the first prompt, a behavior that rarely occurred in subsequent attempts. However, from the third prompt onward, participants appeared to grasp the independence of each attempt, as evidenced by a marked decrease in cross-prompt references and a corresponding rebound in performance.
Next, the bottom pane of Figure 1B shows the average treatment effect (ATE) per iteration, which is the difference between the stratified averages of DALL-E 3 and DALL-E 2 in the top pane. [^4] In Section D, when we refer to “DALL-E 3”, we mean “DALL-E 3 (Verbatim)” unless otherwise specified. To test the widening impact of using DALL-E 3 on performance relative to DALL-E 2, we also run the following fixed effects linear model with participant-level ($i$) clustered standard errors where iteration is treated as a numeric variable:
$Y_{i,n,t}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$ (6)
The coefficient estimates generated by this model are:
  • $\hat{\beta_{1}}=0.0011$, $\hat{SE}(\beta_{1})=0.0003$, $p=0.0004$
  • $\hat{\beta_{2}}=0.0120$, $\hat{SE}(\beta_{2})=0.0037$, $p=0.0013$
  • $\hat{\beta_{3}}=0.0010$, $\hat{SE}(\beta_{3})=0.0004$, $p=0.0227$
The overall ATEs that we report between different pairs of treatment arms (DALL-E 2, DALL-E 3, and DALL-E 3 with revisions) in the main text are estimated from a two-way fixed effect (iteration and target image) model per each pair. Standard errors are cluster robust at the participant level.
D.2 Prompt Characteristics
Figure 1C compares the prompt length and prompt similarity of the two models. To generate these results, we first remove any prompt that does not constitute a good-faith attempt according to the sample construction procedure detailed in Section B.B.4. The prompt length is the average number of words per model and iteration stratified by the reference image. The prompt similarity is the average cosine similarity between all consecutive pairs of user prompts, which are both determined to be valid attempts, stratified by the reference image (see Section B.B.5.3 for details on similarity calculations). The color scale in figure 1C shows the stratified average similarity to the previous prompt across all users per each model. We find that superior model users write prompts that, on average, have $\beta=0.0184$ higher in cosine similarity to their previous prompts using cluster robust standard errors at the participant level ($p=0.0236$).
Comparing the aggregate similarity of all attempts made by a given participant, we also find the prompts from DALL-E 3 participants were, on average, more similar than the prompts of the inferior model participants. For this analysis, we use the dispersion around the centroid in the prompt embedding space, explained in Section B.B.5.3, as the dependent variable. When we average across all participants by model, we find that the average distance of prompts written by superior model participants to their centroid is $\beta=0.0191$ smaller than inferior users ($p=0.0083$). Standard errors are cluster-robust at the participant level.
D.3 ATE Decomposition
Figure 2 in the main text decomposes the ATE into the model and prompting effects. This decomposition is conceptually similar to a simple mediation analysis, with an important difference being that we can observe counterfactual outcomes (e.g., prompting the superior model as if it is the inferior model). This is not typically the case in mediation analysis, and makes causal identification rely on fewer assumptions. To obtain counterfactual outcomes, we fed or “replayed” the participant prompts when interacting with one model (e.g., inferior) on another model (e.g., superior). The notation (prompt, model) specifies which treatment arm the prompts were written under and which model was used in the replay. For example, (2,3) indicates replaying prompts written under DALL-E 2 on DALL-E 3. To be clear, (2,2) and (3,3) correspond to the original observed treatment arms, while (2,3) and (3,2) are the counterfactual outcomes of interest. [^5] To avoid problems with model drift, we regenerated images for all four possible combinations at the same time and used these images for all analyses in the main text.
The left-most point in Figure 2 corresponds to the average CLIP cosine similarity to the target image of (2,2). To make the interpretation of the results clearer, we have subtracted this quantity from all average quality scores and added a dashed line throughout. The second point from the left corresponds to average similarity to the target of (2,3), the third point from the left to (3,3), and the rightmost points to (3,2). All average similarity scores are stratified by iteration and reference image, and the standard errors are bootstrapped and cluster-robust at the participant level. The model effect, as shown by the red braces in Figure 2, corresponds to the average increase in quality of (2,3) relative to (2,2). In the terminology of mediation analysis, the model effect would be referred to as the direct effect. The prompting effect, as shown by the blue braces in Figure 2, corresponds to the average increase in quality of (2,3) relative to (3,3). In the terminology of mediation analysis, the prompting effect would be referred to as the indirect effect. We can also test the difference in average quality between (3,3) and (3,2), as well as the difference between (3,2) and (2,2). Both of these differences are visible in Figure 2; the second is small and not statistically significant.
The standard errors in Figure 2 correspond to the uncertainty around the estimated average score for each of the four replay conditions. These uncertainty estimates are insufficient for exact inference on the direct, indirect, and treatment effects. The statistics and significance values report in the main text, which correspond to such effects (i.e., the difference between average estimates in two conditions) are obtained using a two-way (iteration and target image) fixed effect model with the effect type as the main independent variable:
$Y_{i,n,t}=\beta_{0}+\beta_{1}\text{effect}+\alpha_{n}+\gamma_{t}+\epsilon_{i,n,t}$ (7)
where $\beta_{1}$ is the coefficient on the effect type in question (i.e., model or prompting). To estimate the different effects, we simply use the above model and filter the data as appropriate. For example, to estimate the ATE, the data contains all (2,2) and (3,3) scores, and in this case effect=1 for observations in (3,3) group. Similarly, to estimate the direct or model effect, the data contains all (2,2) and (2,3) scores and effect=1 for observations in (2,3) group. Finally, to estimate the indirect or prompting effect, the data contains all (2,3) and (3,3) scores and effect=1 for observations in (3,3) group. The standard errors for each estimated model are cluster robust at the participant level, and p-values are adjusted accordingly.
Appendix E Robustness Checks
E.1 DreamSim-based Analysis
As discussed in Section B, we repeat all main-text analyses with DreamSim. The results of these analyses are reported below.
E.1.1 Overall ATEs
In terms of DreamSim, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.238$ standard deviations (95% CI = [0.152, 0.324]) closer to the target image ($\Delta DreamSim=0.0306$, $p<10^{-7}$) than those produced by participants using DALL-E 2 (the inferior model).
E.1.2 Figure 1
We reran the regression in Section D.D.1 with $Y_{i,n,t}$ representing the DreamSim outcome instead of cosine similarity:
$Y_{i,n,t}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\gamma_{t}+\epsilon_{i,n,t}$ (8)
The coefficient estimates generated by this analysis are:
  • $\hat{\beta_{1}}=0.0034$, $\hat{SE}(\beta_{1})=0.0005$, $p=1.3\times 10^{-12}$
  • $\hat{\beta_{2}}=0.0200$, $\hat{SE}(\beta_{2})=0.0061$, $p=0.0011$
  • $\hat{\beta_{3}}=0.0024$, $\hat{SE}(\beta_{3})=0.0007$, $p=0.0009$
E.1.3 Figure 2
Decomposing the ATE as measured in terms of DreamSim, we find similar results to those in the main text. The model effect accounts for 54.4% of the ATE ($\Delta DreamSim=0.0166$, $p<10^{-7}$), whereas the prompting effect accounts for 45.4% of the ATE ($\Delta DreamSim=0.01390$, $p=0.014$).
E.2 Z-score-based Analysis
As we discuss in Section C.2, we repeat all main-text analyses with the within-image-attempt Z-score of CLIP-cosine similarity to better account for variation between images and attempts. When comparing across attempts, the Z-score is computed within each image, as mentioned in section C.2. The results hold across the board and sometimes, the differences between the superior and inferior model are even starker than when using cosine similarity.
E.2.1 Overall ATEs
As mentioned in the main text, participants using DALL-E 3 (the superior model) produced images that were, on average, $z=0.19$ standard deviations (obtained from ATE in terms of Z-Scored Cosine Sim = 0.19, 95% CI = [0.100, 0.271]) closer to the target image ($\Delta CoSim=0.0164$, $p<10^{-5}$) than those produced by participants using DALL-E 2 (the inferior model). Standard errors are clustered at the participant level.
E.2.2 Figure 1
On average, participants using the superior model produced images that were $z=0.19$ standard deviations closer (the ATE) to the target image than those using the inferior model. Like in the main text with CLIP cosine similarity, this treatment effect increased as participants made successive attempts to replicate the target image.
$\text{ZScore}_{i,n}=\beta_{0}+\beta_{1}\text{iteration}+\beta_{2}\mathbb{I}[\text{dalleVersion = 3}]_{i}+\beta_{3}\text{iteration}\times\mathbb{I}[\text{dalleVersion = 3}]_{i}+\epsilon_{i,n}$ (9)
  • $\hat{\beta_{1}}=0.0129$, $\hat{SE}(\beta_{1})=0.0038$, $p=0.0007$
  • $\hat{\beta_{2}}=0.1250$, $\hat{SE}(\beta_{2})=0.0457$, $p=0.0064$
  • $\hat{\beta_{3}}=0.0128$, $\hat{SE}(\beta_{3})=0.0053$, $p=0.015$
E.2.3 Figure 2
When we decompose the ATE into the model effect ($z=0.0791$; $p=8.35\times 10^{-6}$) and prompting effect ($z=0.1046$; $p=0.016$), they account for 43% and 56% of the treatment effect, respectively. And when we replay the inferior model prompts on the superior model, the difference in similarity to the target from these prompts played on the inferior is not statistically significant and close together ($z=-0.033$; $p=0.45$). In short, Figure 2 in the main text is quantitatively and qualitatively unchanged when using the within-image Z-score of the cosine similarity.
Appendix F Pre-registration
Prior to our experiment, we pre-registered a number of hypotheses and a pre-analysis plan. This pre-registration is deposited at OSF at the following URL: https://osf.io/ejbtp. The main text of this paper contains a subset of our pre-registered analysis, as well as complementary exploratory analyses that are also mentioned in our pre-registration. We chose to present this subset of our pre-registered analyses because we believe this subset constitutes an important and timely set of results best-suited to a short-form paper.
Below, we provide a high-level description of all of our pre-pregistered analyses, along with a description of the results of those analyses. In our pre-registration, we declared the intent to conduct each of our analyses using six666see section F.2 on deviations from the pre-registration for the remaining two pre-registered outcome variables. possible outcome variables, all of which are different transformations of the same underlying data: CLIP embedding cosine similarity and DreamSim, both of which rescaled in three ways:
  • No rescaling: The outcome variable is used as-is.
  • Z-score rescaling: We rescale the outcome variable into a Z-score according to the procedure describe in these supplementary materials.
  • Percentile rank rescaling: We rescale the outcome variable into a percentile rank. This is done by calculating the percentile rank of a given prompt relative to all other prompts submitted for the relevant target image.
We intend to complete each pre-registered analysis with each of these eight possible outcome variables. Although we have not yet done so, we have no reason to believe that our results will not be robust to these different transformations of our data. However, we did not want to delay the publication of our preprint until all of these analyses were complete. Thus, we are posting a preprint with our pre-registered analyses complete for only a subset of these pre-registered outcomes. As we complete more of these analyses, we will update this supplement. We also intend to post full results, data, and replication code online shortly.
F.1 Hypotheses and Results
Below, we list the hypotheses precisely as they are written in our pre-registration document. For each hypothesis, we also describe our results.
H1
There are differences in prompt engineering ability (as measured through metrics such as average expected prompt quality, initial expected prompt quality, and max expected prompt quality) across demographic attributes and other observables, such as educational background and occupational skills.
Analysis approach:
For this hypothesis, we conducted multiple ANOVA tests, one for each of the demographic variables, against the relevant outcome variable and treatment variable (the model) as the covariate. As a robustness check, we repeated the same procedure using the Kruskal-Wallis U test. To adjust for multiple testing, we used the Benjamini-Hochberg adjustment with a false discovery rate of 0.05.
Results: The results of our analysis are as follows:
  • CLIP embedding cosine similarity
    No rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills of critical thinking, active listening and quality control.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking as job skills, and report little usage of generative AI or imagery writing skill are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those with self-reported occupational skill of quality control, and those who programmed more frequently performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables. In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
    Z-score rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability and frequency, outlook towards generative AI, age, gender, generative AI use, education, imagery writing skill, and self-reported occupational skills of critical thinking, quality control, technology design, social perceptiveness and troubleshooting.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and troubleshooting as job skills, and report little usage of generative AI, little imagery writing skill, some programming skill, and some instruction writing skill are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those with a graduate degree, those with self-reported occupational skill of quality control and technology design, and those who programmed more frequently performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables. In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
    Percentile rank rescaling: Among all the demographic variables, the following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability and frequency, outlook towards generative AI, age, gender, generative AI use, and self-reported occupational skill of critical thinking.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking as a job skill, and report little usage of generative AI, and some programming skill are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, we found fewer significant variables. In this case, the statistically significant relationships with performance are: self-reported programming frequency, outlook towards generative AI, self-reported programming skill, gender and age.
  • DreamSim
    No rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery and instructional writing skill, and self-reported occupational skills of critical thinking, learning strategies, technology design and quality control.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, and those who programmed more frequently performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies occupational skill.
    Z-score rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery and instructional writing skills, education, and self-reported occupational skills of critical thinking, social perceptiveness, learning strategies, technology design and quality control.
    When we fit a linear model to evaluate directionality, we find that those who use critical thinking and social perceptiveness as job skills, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently, those with graduate degrees, and those with technology design, quality control and learning strategies as occupational skills performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies, social perceptiveness and technology design occupational skills.
    Percentile rank rescaling: The following variables have a significant association with the similarity to the target image when testing with ANOVA: computer programming frequency, self-reported programming ability, outlook towards generative AI, age, gender, generative AI use, imagery writing skill, education, and self-reported occupational skills of social perceptiveness, learning strategies, and technology design.
    When we fit a linear model to evaluate directionality, we find that those who use social perceptiveness as a job skill, and report little usage of generative AI, report little imagery writing or programming skills are, on average, better at the task. Conversely, we find that older people, men, those with a more positive outlook regarding generative AI, those who programmed more frequently, and those with technology design and learning strategies as occupational skills performed worse on our experiment task.
    When conducting the non-parametric Kruskal-Wallis test, the following variables had a statistically significant relationships with performance: self-reported programming frequency, generative AI use, outlook towards generative AI, self-reported imagery writing skill, gender, age and learning strategies, social perceptiveness and technology design occupational skills.
H2
There are observable differences in the prompting techniques of successful prompt engineers and unsuccessful prompt engineers. Such prompting techniques might include the use of longer prompts, the use of structured prompting techniques, and/or specific patterns in the way that the participant iterates on their prompts over time.
Analysis approach:
To investigate this hypothesis, we pre-registered evaluating how exploration/exploitation in the prompting space is related to performance. Hence, most of the pre-registered independent variables measured the level of similarity (i.e., exploitation) across a user’s prompts. We refer to participants being more “exploitative” if they wrote prompts that are more similar to their previous prompts and more “exploratory” if they wrote prompts that deviated more from their previous prompts. As described below, we operationalized “similarity” in a variety of ways.
  1. 1.
    Positively associated with exploitation: average token sort ratio compared to the previous prompt, average cosine similarity between the embeddings of a given prompt with the previous prompt, fraction of times a prompt contains the previous prompt as an exact substring
  2. 2.
    Negatively association with exploitation: the variance of the prompt embedding, the number of topical transitions
We also pre-registered measuring each participant’s average prompt length. These variables were measured per each user and across their first 10 attempts. Their association with performance, measured in terms of each of our eight outcome variables, was estimated in a linear model with DALL-E version fixed effects.
We also conducted a similar analysis at the iteration level to answer the following question: how is exploration/exploitation associated with subsequent performance in the next attempt, and is the strength of this relationship mediated by the quality of the previous attempt? To answer this question, first, we divided user-iteration observations into 6 equal-sized brackets by performance in the previous iteration. The bracketing allows us to explore heterogeneity in prompting behavior by the quality of previous attempts. We then estimated the effect of textual similarity to the previous prompt on the quality of the next attempt within each bracket. Our estimates adjusted for other covariates by matching user-iteration observations on the target image, DALL-E model, iteration and the exact quality of the previous attempt (Sävje et al., 2021).
Results: The results of our analysis are as follows:
  • CLIP embedding cosine similarity
    No rescaling: At the user-level, we find strong and statistically significant evidence of an association between performance and our prompting variables, even after adjusting for Benjamini-Hochberg multiple testing. Token sort ratio, cosine similarity with the previous prompt, and frequency of including the previous prompt were positively associated with performance, while embedding variance and the number of topical transitions negatively correlated with performance. Taken together, these findings suggest that more successful users engaged in more exploitation and wrote prompts that were similar to one another. We also found that longer prompts were associated with higher performance, as shown in the main text.
    At the user-prompt level, we find that when the previous performance was poor, higher exploration (or lower cosine similarity with the previous prompt) was associated with improved performance, although extremely high levels of exploration did not improve performance. In contrast, in cases where previous performance was high, higher exploitation monotonically increased performance. Using token sort ratio with the previous prompt as the measure of exploration also generated similar results. However, when using binary measures of exploration, i.e. topical transition or containing the previous prompt, we found that exploitation is associated with higher performance in the next iteration regardless of the bracket. As these are binary measures of exploration, we could not replicate the non-linearity we observed with continuous measures of exploration in the low-performing group. However, we do find that topical transitions (more exploration) leads to an overall drop in performance and the drop becomes larger for higher performance in the previous attempt. We find a similar pattern when measuring exploitation by whether a prompt contains the previous one. Performance is overall higher if the prompt includes the previous one and the improvement is larger as the previous performance gets higher.
    Z-score rescaling: We find the same results as those described for the unscaled cosine similarity above. The main difference is that the non-linearity between performance and continuous measures of exploitation, i.e. TSR ratio and cosine similarity with previous prompt, observed at the user-prompt level becomes starker for the bottom two brackets of previous performance. For low-performing prompts in the previous attempt, the optimal exploitation level is in the middle and the performance greatly deteriorates for high and low levels of exploitation.
    Percentile rank rescaling: We find the same results as described above, however the non-linearity in exploration/exploitation vs performance in the bottom two brackets of previous performance is now even stronger than those observed with Z-score rescaled cosine similarity as the performance measure.
  • DreamSim
    No rescaling: We find exactly the same results as those described for the unscaled cosine similarity above.
    Z-score rescaling: We find the same results as those described for the Z-score cosine similarity above. In particular, we again find that the non-linearity between performance and continuous measures of exploitation observed at the user-prompt level is stark for the bottom two brackets of previous performance.
    Percentile rank rescaling: We find the same results as described above for cosine similarity percentile rank rescaling.
H3
There are differences in prompt engineering techniques (as measured through metrics such as prompt length and iteration-to-iteration token sort ratio) across demographic attributes and other observables, such as educational background and occupational skills.
Analysis approach:
To test this hypothesis, we estimated a linear model per each demographic trait as the independent variable, treatment arm fixed effects, and the prompting behaviors outlined below and described in-depth in Section B.B.5 as the outcome variables of interest. To account for multiple testing, we adjusted the p-values for all these models according to the Benjamini-Hochberg procedure.
Results: The results of our analysis are as follows. As a reference, the definitions of these dependent variables are provided in Section B.B.5.
  • Prompt embedding variance: We did not find any significant differences across various demographic traits.
  • Strategic Shifts: We observed statistically significant differences based on age, reported programming and instructional writing frequencies. Younger participants, those with low programming frequency and those with some instructional writing frequency demonstrated decreased topical transitions across their prompts.
  • Successive Prompt Token Sort Ratio: We found significant differences by age, education, programming frequency, and imagery/instructional writing frequencies. Older users, those with post-graduate degrees, those with high programming frequency and writing frequency, both precise instructions and imagery, write prompts that are less similar to each other on average.
  • Successive similarity: We found significant differences by age, gender, education, generative AI outlook, programming skill/frequency, instructional/imagery writing frequency and some occupational skills. On average, older users, males, those with post-graduate degrees, those who reported frequently computer programming, those who reported being strong computer programmers, those with a positive outlook on Generative AI, those who frequently use generative AI, those who report frequently writing instructions or imagery, and those with self-reported occupational skills of critical thinking and social perceptiveness write prompts that are less similar to each other (as measured by cosine similarity) on average.
  • Successive Prompt ‘Contains Previous Prompt’ Dummy: We found significant differences by age, gender, outlook towards generative AI, and imagery writing frequency. Older users and those with neutral outlook toward generative AI are less likely to write prompts that exactly contain the previous. In contrast, males and those with some imagery writing skills are more likely to keep using their previous prompt.
H4
Insofar as the output returned by a generative AI model in response to a prompt is stochastic, the subsequent prompting strategies and prompting outcomes of participants that get lower-than-expected, higher-than-expected, or approximately expected outputs in response to their first prompt are different.
Analysis approach: To test this hypothesis, we first calculated the Z-score of each participant’s realized image observed during our experiment relative to the sampling distribution that we approximated for the prompt that generated that image according to the procedure outlined in Section B.B.3 (see Section C for details on calculating the Z-score). This Z-score quantifies the random variation in observed image quality relative to the prompt’s true underlying quality. Images with higher Z-scores represent instances where the realized image quality exceeded expectations based on the prompt, while lower Z-scores indicate instances where the image quality was randomly lower than expected for a given prompt. By examining the relationship between these Z-scores and subsequent performance and prompting behavior, we can assess the causal impact of this stochasticity on user behavior.
To estimate the relationship between this stochasticity on subsequent prompting behavior and performance, we transformed the Z-score into a trichotomous variable where any Z-score less than -0.45 as “lower-than-expected,” between -0.45 and 0.45 as “expected” and greater than 0.45 as “higher-than-expected.” We then perform two-sample t-tests with comparing the three groups’ relative performance. We also estimate a linear model as a robustness check, regressing participant performance on the trichotomous Z-score variable with treatment arm fixed effects. We also repeat this analysis, treating the Z-score as a continuous variable without the trichotomous transformation. Finally, we also perform this analysis at the user level with the trichotomous variable, where only the Z-score of the observed image of the first prompt is measured, and we test if that affects the average performance of all subsequent user attempts.
Results: The results of our analysis are as follows, with each outcome variable clearly labeled.
  • CLIP embedding cosine similarity
    No rescaling: We find statistically significant evidence that increases in the Z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt. When we test differences in the trichotomous variable, the difference between the top bracket and the bottom bracket in the realized Z-score is statistically significant, with the top bracket having a higher performance in the next attempt than the bottom bracket. But we do not find significant differences between the bottom and middle brackets and between the top and middle brackets. When we estimate a linear model, we do not find a statistically significant relationship between the observed Z-score of the prompt and performance in the next iteration. This is due to the non-linearity of the relationship between these variables that exists separately in the negative and positive range of realized Z-score. Additionally, when we estimate the effect at the user-level, we find differences between the top and middle brackets of the first prompt realized Z-score. However, these differences are barely significant ($p=0.048$) and do not account for multiple testing. When fitting a linear model at the user-level, we do not find a a statistically significant relationship between the observed Z-score of the first prompt and average performance in the subsequent attempts.
    Z-score rescaling: The results are very similar to those described for the unscaled cosine similarity above. In particular, we find a statistically significant difference between the top bracket and the bottom bracket, with the top bracket having a higher performance in the next attempt than the bottom bracket. We don’t find statistically significant effects when comparing the middle with top or bottom brackets, or when fitting a linear model. In contrast to the unscaled measure, we do not find any statistically significant effects at the user-level, neither when comparing the brackets on the realized Z-score of the first image nor when fitting a linear model between the realized quality of the first prompt and the performance in subsequent attempts.
    Percentile rank rescaling: We find the exact same results as those explained above for Z-score cosine similarity.
  • DreamSim
    No rescaling: We find similar evidence that increases in the Z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt. The difference between the top bracket and the bottom/middle brackets in the realized Z-score is statistically significant, with the top bracket having a higher performance in the next attempt than both the bottom and middle brackets. But we do not find significant differences between the bottom and middle brackets. When we estimate a linear model, we do not find a statistically significant relationship between the observed Z-score of the prompt and performance in the next iteration, likely due to the non-linearity that we also observed between performance measured as cosine similarity and realized Z-score of the previous prompt. When we estimate the effect at the user-level, we do not find any statistically significant relationship between realized quality of the first prompt and performance in subsequent attempts, either using the trichotomous variable on Z-score or using a linear model.
    Z-score rescaling: The results are very similar to those described for the unscaled dreamsim score above. In particular, we find a statistically significant difference between the top bracket and both the bottom and middle brackets, with the top bracket having a higher performance in the next attempt. We don’t find statistically significant effects when comparing the middle with the bottom bracket, or when fitting a linear model. Similar to the unscaled measure, we do not find any statistically significant effects at the user-level, neither when using the trichotomous variable nor when fitting a linear model between the realized quality of the first prompt and the performance in subsequent attempts.
    Percentile rank rescaling: We find the exact same results as those explained above for Z-score dreamsim similarity, with the only difference that at the user-attempt level, we observe a statistically significant difference only between the top and bottom brackets. The top bracket in realized quality no longer has a statistically significant improvement in the performance of the next attempt when compared to the middle bracket.
  • Prompting Behaviors
    Prompt Length: e find statistically significant evidence that increases in the Z-score of a realized image causes the subsequent prompt to be longer (either in terms of words or characters). When we test the differences using the trichotomous variable, the difference between the top Z-score bracket and the bottom bracket is statistically significant, with on average longer prompts in the top bracket. However, we do not observe statistically significant differences between either the top or bottom brackets with the middle bracket. When we estimate a linear model at the user-attempt level, we find a similar relationship that is both positive and statistically significant. A unit increase in the Z-score realized quality of an image causes the next prompt to be longer by about 0.3 words on average. Finally, when we estimate the effect at the user-level, we do not find the Z-score of the first prompt to have a statistically significant effect on the length of the subsequent prompts, when using the trichotomous variable or when fitting a linear model.
    Successive Similarity: We find statistically significant evidence that increases in the Z-score of a realized image causes an increase in successive similarity of the prompt. When we test the differences using the trichotomous variable, the difference between the top Z-score bracket (better than expected images) and both the bottom (worse than expected images) and middle (about expected images) brackets is statistically significant, with the top bracket leading to higher similarity with the previous prompt. However, we do not observe a statistically significant difference between the bottom bracket with the middle bracket. When we estimate a linear model, we find a similar relationship that is both positive and statistically significant. An increase in the realized quality (Z-score) of an image causes the next prompt to be more similar to the previous one. When we estimate the effect at the user-level, we do not find statistically significant effects of the first prompt realized quality on average successive similarity, neither when using the trichotomous variable nor when fitting a linear model.
    Successive Prompt Token Sort Ratio: The results are identical to the case of successive cosine similarity described above. We find statistically significant evidence that increases in the Z-score of the image realized in a previous prompt causes an increase in the Token Sort Ratio between successive prompts, using both the trichotomous variable (top vs bottom and middel brackets) and the linear model. At the user-level, we do not find statistically significant effects of the Z-score of the first prompt on average token sort ratio of successive prompts.
    Successive Prompt ‘Contains Previous Prompt’ Dummy: We find statistically significant evidence that increases in the Z-score of a realized image causes an increase in the likelihood that the subsequent prompt contains the previous one. When we test the differences using the trichotomous variable, the difference between the top Z-score bracket and the bottom bracket is statistically significant, with the top bracket more likely to exactly include the previous prompt. However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket. When we estimate a linear model at the user-attempt level, we find a similar relationship that is both positive and statistically significant. An increase in the realized quality of an image causes the next prompt more likely to include the previous one. Finally, when we estimate the effect at the user-level, we do not find the Z-score of the first prompt to have a statistically significant effect on the average probability of a prompt containing the preceding prompt, when using the trichotomous variable or when fitting a linear model.
H5
Average prompt engineering ability (as measured through metrics such as average expected prompt quality, initial expected prompt quality, and max expected prompt quality) and prompting strategies will depend on the capacity of the model that participants are interacting with.
Analysis approach:
To test this hypothesis, we conducted three two-sample t-tests that compare the prompting performance and prompting behavior of the three possible pairs of treatment assignments. All outcome variables are specified in the results below. For robustness, we repeat these analyses with ANOVA to test whether any of the three treatment arms has an effect on the same variables. To account for multiple testing, we adjusted the p-values for tests with the Benjamini-Hochberg procedure.
Results: The results of our analysis are as follows:
  • CLIP embedding cosine similarity
    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores, as opposed to task-iteration Z-scores). When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks). When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
  • DreamSim
    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores). When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find the same pattern of head-to-head statistical significance results. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks). When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
  • Prompting Behaviors
    Mean prompt Length: We found statistically significant differences in prompt length between treatment arms. Participants using DALL-E 2 used significantly shorter prompts compared to both DALL-E 3 (Revised) and DALL-E 3 (Verbatim) groups. There was no significant difference in prompt length between the two DALL-E 3 groups. ANOVA results confirmed a significant effect of DALL-E version on prompt length.
    Aggregate Similarity: We found no statistically significant differences between any of the treatment arms. This suggests that the overall variability in prompts was similar across all three DALL-E versions. ANOVA results confirmed no significant effect of DALL-E version on this measure.
    Successive Similarity: Analysis of the average cosine similarity between successive prompts revealed no statistically significant differences between the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on successive prompt similarity.
    Successive Prompt Token Sort Ratio: We found no statistically significant differences between the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on this measure.
    Successive Prompt ”Contains Previous Prompt” Dummy: Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences between any of the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on this measure.
H6
Variability in participants’ ability to prompt engineer effectively and prompting strategies will depend on the capacity of the model that participants are interacting with.
Analysis approach:
To test this hypothesis, we conducted two analyses. First, would conducted F-tests comparing the variance of participant performance and prompting behaviors between all 3 pairs of models. Second, we estimated the quantile treatment effects (QTEs) between all 3 pairs models on participant performance and prompting behaviors. We also pre-registered our intent to visually inspect whether the QTEs we observe are consistent with dispersion/“inequality” being reduced or increased when participants use different models (e.g., positive effects for low quantiles and negative/null effects for high quantiles would be consistent with inequality reduction).
F-test Results (with BH Adjusted p-values):
  • DALL-E 3 (revised) vs. DALL-E 2
    Mean prompt length (words): DALL-E 3 Revised has significantly less variance than DALL-E 2 (ratio 0.5217, $p\leq 10^{-4}$).
    Prompt embedding variance: DALL-E 3 Revised also has significantly less variance than DALL-E 2 (ratio 0.7123, $p=2\times 10^{-4}$).
    Cosine similarity with target image, Z-Score: DALL-E 3 Revised has significantly more variance than DALL-E 2 (ratio 1.255, $p=0.0159$).
    Failed to reject null of no differences in variance: Mean raw DreamSim score vs. target image (p = 0.1333), number of topical transitions (by token sort ratio) (p = 0.3294), mean token sort ratio with previous prompt (p = 0.6244), mean percentile rank of DreamSim score (p = 0.9035), mean cosine similarity to previous prompt (p = 0.8748), mean raw CosineSim score vs. target image (p = 0.8301), mean proportion of prompts containing previous prompt (p = 0.5461), Z-score of DreamSim score (p = 0.4016), number of topical transitions (cosine similarity) (p = 0.3505), and mean percentile rank of CosineSim (p = 0.0795).
  • DALL-E 3 (verbatim) vs. DALL-E 3 (revised)
    Mean percentile rank of CosineSim: DALL-E 3 (verbatim) has significantly less variance than DALL-E 3 (revised) (ratio 0.7347, $p=0.0009$).
    Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) also has significantly less variance (ratio 0.7352, $p=0.0009$).
    Mean percentile rank of DreamSim: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7888, $p=0.0141$).
    Cosine similarity with target image, Z-Score: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7961, $p=0.0159$).
    DreamSim with target image, Z-Score: DALL-E 3 (verbatim) has significantly less variance (ratio 0.8195, $p=0.0377$).
    Number of topical transitions (token sort ratio): DALL-E 3 (verbatim) has significantly more variance (ratio 1.2303, $p=0.0301$).
    Failed to reject null of no differences in variance: Mean raw DreamSim score vs. target image (p = 0.1842), mean raw CosineSim score vs. target image (p = 0.5916), prompt embedding variance (p = 0.9771), mean cosine similarity to previous prompt (p = 0.9035), mean prompt length (words) (p = 0.6244), number of topical transitions (cosine similarity) (p = 0.5916), and mean token sort ratio with previous prompt (p = 0.5206).
  • DALL-E 2 vs. DALL-E 3 (verbatim)
    Mean prompt length (words): DALL-E 3 (verbatim) has significantly less variance than DALL-E 2 (ratio 0.553, $p\leq 10^{-4}$).
    Prompt embedding variance: DALL-E 3 (verbatim) also has significantly less variance (ratio 0.7157, $p=2\times 10^{-4}$).
    Mean raw DreamSim score vs. target image: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7501, $p=0.0015$).
    Mean proportion of prompts containing previous prompt: DALL-E 3 (verbatim) has significantly less variance (ratio 0.7904, $p=0.0141$).
    Mean percentile rank of DreamSim: DALL-E 3 (verbatim) has significantly less variance (ratio 0.8005, $p=0.0159$).
    Failed to reject null of no differences in variance: Mean percentile rank of CosineSim (p = 0.1842), Z-score of DreamSim score (p = 0.3294), Mean Raw CosineSim (p = 0.8151), Z-score of CosineSim score with target image (p = 0.9906), mean token sort ratio with previous prompt (p = 0.8748), mean cosine similarity to previous prompt (p = 0.807), number of topical transitions (token sort ratio) (p = 0.3505), and number of topical transitions (cosine similarity) (p = 0.085).
QTE Highlighted Results: Through our QTE plots, we see clearest evidence of dispersion being reduced for:
  • Z-score of cosine similarity with respect to the target image (calculated within task-iteration) for the DALLE-3 Revised vs. DALLE-2 comparison.
And we see the clearest evidence of dispersion increasing for:
  • Mean prompt length for all three pairwise model comparisons (DALLE-3 Verbatim vs. DALLE-2, DALLE-3 Revised vs. DALLE-2, and DALLE-3 Verbatim vs. DALLE-3 Revised)
  • Raw Dreamsim performance for DALLE-3 Verbatim vs. DALLE-2 and DALLE-3 Revised vs. DALLE-2 comparisons.
H7
As participants repeatedly try to complete a task with a given model, the quality of their attempts will increase, and the extent to which the quality increases varies as a function of model capacity.
Analysis approach:
In order to test for differences in prompting performance improvement across participant iterations, we pre-registered and conducted stratified two-sample tests with the 3 treatment arms as our strata. The two samples represent the best and initial scores. We make these comparisons overall and within strata.
  • CLIP embedding cosine similarity
    No rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
    Z-score rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
    Percentile rank rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
  • DreamSim
    No rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
    Z-score rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
    Percentile rank rescaling: For both within strata and overall, we see the best scores outperforming the initial scores, in a statistically significant way.
H8
The extent to which participants can recreate images using models such as DALL-E 2/3 will vary across images.
Analysis approach:
To evaluate this hypothesis, we pre-registered and performed two analyses. First, we compared the extent to which different images can be replicated using GPT-4V—a multimodal generative AI model that can take in both text and images and output text. For each of the 15 target images, we prompted GPT-4V to “Write a DALL-E {2, 3} prompt to recreate this image verbatim as closely and as detailed as possible.” We do this once for DALL-E 2 and once for DALL-E 3, generating two “AI prompts” per target image. We then provided these AI prompts to DALL-E 2, DALL-E 3 (Verbatim), and DALL-E 3 (Revised), respectively, 20 times each, generating 60 replicated images per target image (the DALL-E 3 AI prompt was sent to both DALL-E 3 (Verbatim) and (revised)). For each image, we measure the cosine similarity of the CLIP embedding vectors to those of the relevant target image. We then average these 60 similarity measures by target image, resulting in a mean similarity score that represents GPT-4V’s ability to generate prompts that recreate each target image. Second, we simply measured the similarity of every participant-generated image to that of the relevant target image, and averaged over these participant-generated similarities.
Results:
  • CLIP embedding cosine similarity: Below is the ranking of GPT-4V’s ability to generate prompts that recreate each target image, as measured through CLIP embedding cosine similarity:
    1. 1.
      Business Image #3
    2. 2.
      Business Image #5
    3. 3.
      Business Image #1
    4. 4.
      Photography Image #1
    5. 5.
      Design Image #2
    6. 6.
      Business Image #2
    7. 7.
      Photography Image #5
    8. 8.
      Design Image #4
    9. 9.
      Business Image #4
    10. 10.
      Design Image #1
    11. 11.
      Photography Image #2
    12. 12.
      Design Image #5
    13. 13.
      Design Image #3
    14. 14.
      Photography Image #3
    15. 15.
      Photography Image #4
    The maximum average cosine similarity of the CLIP embeddings in the above list (the easiest image for GPT-4V to replicate) was $CoSim=0.944$ for Business Image #3, and the lowest score (the hardest image for GPT-4V to replicate was $CoSim=0.734$ for Photography Image #4.
    Below is the ranking of participants’ ability to generate prompts that recreate each target image, as measured through CLIP embedding cosine similarity:
    1. 1.
      Business Image #3
    2. 2.
      Business Image #5
    3. 3.
      Business Image #1
    4. 4.
      Photography Image #5
    5. 5.
      Design Image #2
    6. 6.
      Design Image #4
    7. 7.
      Business Image #2
    8. 8.
      Design Image #5
    9. 9.
      Design Image #1
    10. 10.
      Photography Image #1
    11. 11.
      Design Image #3
    12. 12.
      Photography Image #2
    13. 13.
      Photography Image #3
    14. 14.
      Photography Image #4
    15. 15.
      Business Image #4
    The maximum average cosine similarity of the CLIP embeddings in the above list (the easiest image for participants to replicate) was $CoSim=0.892$ for Business Image #3, and the lowest score (the hardest image for participants to replicate was $CoSim=0.669$ for Business Image #4.
  • DreamSim: Below is the ranking of GPT-4V’s ability to generate prompts that recreate each target image, as measured through 1-DreamSim:
    1. 1.
      Design Image #3
    2. 2.
      Business Image #2
    3. 3.
      Business Image #4
    4. 4.
      Design Image #2
    5. 5.
      Business Image #1
    6. 6.
      Business Image #5
    7. 7.
      Design Image #4
    8. 8.
      Business Image #3
    9. 9.
      Photography Image #1
    10. 10.
      Photography Image #5
    11. 11.
      Photography Image #2
    12. 12.
      Design Image #1
    13. 13.
      Design Image #5
    14. 14.
      Photography Image #3
    15. 15.
      Photography Image #4
    The maximum 1-DreamSim score in the above list (the easiest image for GPT-4V to replicate) was $\tilde{D}=0.75$ for Design Image #3, and the lowest score (the hardest image for GPT-4V to replicate was $\tilde{D}=0.40$ for Photography Image #4.
    Below is the ranking of participants’ ability to generate prompts that recreate each target image, as measured through 1-DreamSim:
    1. 1.
      Business Image #3
    2. 2.
      Business Image #2
    3. 3.
      Business Image #5
    4. 4.
      Business Image #1
    5. 5.
      Design Image #2
    6. 6.
      Design Image #3
    7. 7.
      Photography Image #2
    8. 8.
      Photography Image #5
    9. 9.
      Design Image #5
    10. 10.
      Design Image #4
    11. 11.
      Business Image #4
    12. 12.
      Photography Image #1
    13. 13.
      Photography Image #4
    14. 14.
      Design Image #1
    15. 15.
      Photography Image #3
    The maximum 1-DreamSim score in the above list (the easiest image for participants to replicate) was $\tilde{D}=0.575$ for Design Image #3, and the lowest score (the hardest image for participants to replicate was $\tilde{D}=0.356$ for Photography Image #4.
Exploratory Analyses
In our pre-registration, we also mention a number of exploratory analyses that are not described with the same level of detail. Multiple of these exploratory analyses appear in our main text. These pre-registered exploratory analyses are copied verbatim below:
“We plan to investigate whether differences in prompt engineering ability across demographic and other observed variables will vary depending on the complexity of the task, e.g., the difficulty of the image participants are being asked to replicate. We anticipate power for this analysis will be very low, so we chose to label it as an exploratory analysis rather than a pre-registered hypothesis.
We anticipate that we may conduct additional analysis of the prompts submitted by participants (and how these prompts evolve over the course of a session). Furthermore, we might explore the tips that participants provide after completing the task on how to prompt engineer effectively.
We also may take original and revised prompts submitted to DALL-E 3 treatment arms and submit them to DALL-E 2 (and vice versa) to see how participants would have counterfactually performed under different treatment assignments than the one to which they were assigned.”
F.2 Deviations From Pre-registration
Here, we also report any deviations from our pre-registered analysis. By and large, these deviations occurred because certain aspects of our pre-registration were not appropriate from a statistical analysis perspective, or were infeasible.
  • In our pre-registration, we had anticipated using t-tests and Mann-Whitney U tests for many hypotheses. However, this turned out to be impossible for many variables since most of the demographic traits have multiple categories. Thus, we applied ANOVA and Kruskal-Wallis tests instead.
  • In our pre-registration, we were not explicit enough on how to compute Z-score measures of performance. As described in section C.2, Z-scores are almost always computed within image-attempt to adjust for variations across target images or attempts. The only exception is with any analysis that compares performance across attempts where Z-scores need to computed within target images. The pre-registration anticipated that Z-scores for any analysis would be computed only within target images.
  • In our pre-registration, we had forgotten to include Education and GenAI outlook in our list of demographics with respect to which we would measure task performance heterogeneity. We decided to include these important variables in our analysis, despite the accidental omission.
  • In addition to our pre-registered prompt exclusion criteria, we removed additional prompts that did not appear to be “good-faith efforts” to complete our task based on the text of those prompts (see Section B.4 for more details). We find that our results are the robustness to the inclusion of prompts that were removed by this procedure.
  • Because they were easier to implement, we ran t-tests rather than z-tests to conduct our tests of H5. Because the two tests are asymptotically equivalent, we do not believe this will make a difference in our analysis.
  • When conducting our analyses to test H4, we observed that the distribution of the Z-scores (our independent variable) did not conform to a normal distribution, displaying extremely large (7.1) or small (-21.9) values. To prevent our findings from being disproportionately influenced by these outliers, particularly in linear models, we excluded observations with absolute Z-scores greater than 3, which constituted 2.65% of the user-attempt observations. For robustness, we also conducted the stratification analysis including these outliers and found almost identical results, with the exception that the difference in performance between the top bracket and middle or bottom brackets was in some cases only marginally significant.
  • In our pre-registration, we had included a rescaled version of our main performance metrics described in B.5.1 based on GPT-4V for robustness check. This outcome variable would have been rescaled by measuring its distance to the outcome variable obtained using a prompt generated by GPT-4V in response to the target image. Given that the quality of the prompt from GPT-4V model is a constant for each target image, this would have amounted to subtracting a constant from the unscaled CLIP cosine similarity or dream sim scores. Furthermore, since all our analysis involving performance as the dependent variable adjust for the target image, either through post-stratification or as a covariate in a linear model, the results from using GPT-4V scaling would have been identical to the the unscaled measures of performance. For this reason, we have not included the robustness checks involving GPT-4V for our pre-registered hypothesis.